suppose that in a classroom of 30 students 14 are member of the matlab 10 are member of the science club and for our member of both math and science club find the probability that a student will be a member of the month club or student club or both


1
Expert's answer
2021-07-05T17:38:30-0400

If in a classroom of 30 students 14 are members of the matlab, 10 are members of the science club and 4 are members of both math and science clubs, then

only members of matlab club are 14 - 4 = 10,

only members of science club are 10 - 4 = 6, so

10 + 6 + 4 = 20 participate in matlab, science or both clubs.

So, the probability that a student will be a member of the math club or student club or both is

P = 20/30 = 0.67.